OpenStudy (anonymous):

Please help: A toy that cost x$ is sold for $24 for a x% profit. wth is this even possible to solve? (under quadratic problem)

OpenStudy (anonymous):

oh i see a \(x\%\) profit that is harder

OpenStudy (anonymous):

yeah I cna't figure it out... I'm thinking maybe there is missing or incorrect information or something... my brain keeps thinking how can anyone know how much profit and cost of anything sold by only knowing how much it's sold for??? :(

OpenStudy (anonymous):

well maybe we could work with an example if you buy it for $20 and sell if for $24 then you make a \(\frac{4}{20}=.2=20\%\) profit

OpenStudy (anonymous):

damn got it on the first try!! that was just a guess, but it works i see

OpenStudy (anonymous):

lol yeah but you had another number lol

OpenStudy (anonymous):

im supposed to solve it without using 20 ... you are right though the answer it says is 20 hehehe

OpenStudy (anonymous):

i like that answer looks good to me

OpenStudy (anonymous):

ok lets do it without the 20, i was just using that as an example anyway

OpenStudy (anonymous):

lets replace 20 by x

OpenStudy (anonymous):

okay :D

OpenStudy (anonymous):

you can't use 4 either :P

OpenStudy (anonymous):

you buy for \(x\) you sell for \(\$24\) and your profit is \((24-x)\) which, as a percent, is \[\frac{24-x}{x}\times 100\]

OpenStudy (anonymous):

set \[x=\frac{100(24-x)}{x}\] and solve for \(x\)

OpenStudy (anonymous):

hmm that looks good

OpenStudy (anonymous):

you get \[x^2=2400-100x\] \[x^2+100x-2400=0\]

OpenStudy (anonymous):

excellent! :) ill try solving that now thanks

OpenStudy (anonymous):

yeah by some miracle this factors, and of course you know how, since we already know the answer that is why i wanted to do it with numbers, to see how it would work with the variables always good if you get confused

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

yeah that is weird hehe